OpenStudy (honeymonster):

help simplifying this fraction (6x+8x^2)/2x^2

OpenStudy (honeymonster):

\[\frac{ 6x+8x^2 }{ 2x^2 }\]

OpenStudy (anonymous):

factor out 2x

OpenStudy (honeymonster):

how?? \[\frac{ 2x\left( 3+4x\right) }{ 2x }\]??

OpenStudy (swag):

4x^4+3x^3

OpenStudy (anonymous):

(3 +4x)/x

OpenStudy (honeymonster):

so which one is correct?

OpenStudy (anonymous):

SWAG multiply it but its a division since "/"